How many grams of C2H2 will be produced, if 7.00g of Ca(OH)2 are also produced in the following reaction? Show the solution.

1
Expert's answer
2022-05-31T13:52:03-0400

CaC2 + 2H2O = Ca(OH)2 + C2H2

molecular mass of Ca(OH)2 = 74 g/mol

molecular mass of C2H2 = 26 g/mol.

According to the above reaction, when 74 g of Ca(OH)2 is formed, 26 g of C2H2 is formed. When 7 g of Ca(OH)2 is formed, we find how many g of C2H2 are formed:

74 g Ca(OH)2 — 26 g C2H2

7 g Can(OH)2 — X

X = 7 * 26 / 74 = 2,46 g

ANSWER:. 2,46 g C2H2